• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 08:09:31 on June18,2010

    Entry number 327239

    keyword=Owl shift summary

    00:00----Shift starts.


    Time---Up Colli---Down Colli---rad44_p1---rad44_p2---Pressure--Temp
    01:03--0.434------0.406--------0.085------0.007------9e-7------27.2

    01:40----Parity feedback moniotor was not working, bob reduced the gains for 70 uA and PANGUIN was cutting the data for lobeam. Luis called and helpd to parity feedback working. Its working now. We have started teh beam recovery procedure.


    01:54----We did spot++, LHRS run 27904, raster is 5x4 (23x9.5 MCC). I chaned it to 6x4 (27x9.5 MCC).

    02:58----We have raster 6x4 LHRS run 27905.

    03:04----Halog 327172 shows red dots at the left and right side of "Raster X vs Y". I am afraid that may be we are hitting the frame. I changed the raster from 27x9.5 to 23x9.5 which is 6x4 mm to 5x4 mm.

    06:10----Right arm detectors showed low signal. Bob and Jon came and tried to fix it but the problems remains. Bob decided to start production with one arm and go to 70 uA. We went to 70 ua then MCC has some problem and now beam is off for next one hour.